What is the angle between vector p × vector Q and vector Q ×vector P
 

Grade:11

2 Answers

Eshan
5 years ago
Dear student,

The direction of vector obtained by cross product of two vectors is given by the Screw rule. Hence the direction of\vec{P}\times \vec{Q}is opposite to that of\vec{Q}\times \vec{P}. In fact,
\vec{P}\times \vec{Q}=-\vec{Q}\times \vec{P}
Hence the angle between the two vectors is180^{\circ}
Nameet Mishraa
5 years ago
VectorP ×VectorQ -It is the cross product of the vectors p and q and it will lie in a plane perpendicular to that of P and Q.Its direction could be determined by right hand thumb rule..
VectorQ×VectorP -Same like mentioned above.This will lie in plane perpendicular to the vectors P and Q but opposite to Vector P ×Vector Q .
Thus.. since VectorP×VectorQ and VectorQ×VectorP are opposite in direction,So the angle between them would be 180 degrees.